OpenStudy (anonymous):

x^4-13x^2+36=0 Please help. :)

OpenStudy (cwrw238):

hint: let x^2 = X and solve the quadratic

OpenStudy (anonymous):

I get as far as u=9 and u=4 but am unsure what to do after that to get the actual answer.

OpenStudy (cwrw238):

ok u = x^2 i take it so x^2 = 9 and x^2 = 4 so you now can get 4 values of x its an equation in x^4 so there will be 4 roots

OpenStudy (cwrw238):

so 2 roots are sqrt9 , -sqrt9 = -3 and 3

OpenStudy (anonymous):

\[\pm3 and \pm2\]

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

Okay. :)

OpenStudy (cwrw238):

right